OpenStudy (anonymous):

can someone help me find the solution to the equation log16 32 = x + 5 ?

OpenStudy (cwrw238):

log to base 16 right?

OpenStudy (cwrw238):

if so 16^(x+5) = 32

OpenStudy (shubhamsrg):

altough x is simply log(16)32 - 5 but for simplification,, you may break 32 into 16*2..and use log(ab) = log(a+b)

OpenStudy (anonymous):

yeah the log base is 16

OpenStudy (anonymous):

but how would i solve this?

OpenStudy (shubhamsrg):

log(16)32 is a constant term,,why you need to solve it? altough even if you wanna simplify,,as i said,,use 32=16*2 and log(ab)=log a + log b

OpenStudy (cwrw238):

or you can do it this way take logs of both sides: log 16^(x+5) = log 32 by laws of logarithms log a^x = x log a: (x+5) log 16 = log 32 x+5 = log 32/log 16 = 1.25 x = 1.25 - 5 = -3.75

OpenStudy (shubhamsrg):

well you changed bases!! smart buddy!! ;)

OpenStudy (anonymous):

am i dividing here? x+5 = log 32/log 16 = 1.25